A car makes a displacement of 100m towards east and 200m towards north.find the magnitude and direction of the resultant explain deeply

Grade:11

1 Answers

Arun
5 years ago
Dear Madhan
 
 
resultant displacement is (100+2002)1/2 =(10000+40000)1/2=500001/2=100(5)½
the direction is North east
 
Regards
Arun (askIITians forum expert)